Chapter 2

Poisson processes

[28/96%]
Show
LaTeX
Problem 2.1
?
(a)

Find the Erlang density fSn(t)f_{S_n}(t) by convolving fX(x)=λexp(−λx)f_X(x) = \lambda \mathrm{exp}\left(-\lambda x\right) with itself nn times.

(b)

Find the moment generating function (MGF) of XX (or find the Laplace transform of fX(x)f_X(x)), and use this to find the moment generating function (or Laplace transform) of Sn=X1+X2+⋯+XnS_n = X_1+X_2+\cdots +X_n. Invert your result to find fSn(t)f_{S_n}(t).

(c)

Find the Erlang density by starting with (2.15) and then calculating the marginal density for SnS_n.

Problem 2.2
?
(a)

Find the mean, variance, and MGF of N(t)N(t), as given by (2.17).

(b)

Show by discrete convolution that the sum of two independent Poisson rv's is again Poisson.

(c)

Show by using the properties of the Poisson process that the sum of two independent Poisson rv's must be Poisson.

Problem 2.3

The purpose of this exercise is to give an alternative derivation of the Poisson distribution for N(t)N(t), the number of arrivals in a Poisson process up to time tt. Let λ\lambda be the rate of the process.

?
(a)

Find the conditional probability P(N(t)=n∣Sn=τ)\mathbb {P}\left(N(t) = n \mid S_n = \tau \right) for all τ≤t\tau \leq t.

(b)

Using the Erlang density for SnS_n, use (a) to find P(N(t)=n)\mathbb {P}\left(N(t) = n\right).

Problem 2.4

Assume that a counting process {N(t);t>0}\left\{ N(t); t>0\right\} has the independent and stationary increment properties and satisfies (2.17) (for all t>0t > 0). Let X1X_1 be the epoch of the first arrival and XnX_n be the interarrival time between the (n−1)(n-1)th and the nnth arrival. Use only these assumptions in doing the following parts of this exercise.

?
(a)

Show that P(X1>x)=e−λx\mathbb {P}\left(X_1 > x\right) = e^{-\lambda x}.

(b)

Let Sn−1S_{n-1} be the epoch of the (n−1)(n-1)th arrival. Show that P(Xn>x∣Sn−1=τ)=e−λx\mathbb {P}\left(X_n > x \mid S_{n-1} = \tau \right) = e^{-\lambda x}.

(c)

For each n>1n > 1, show that P(Xn>x)=e−λx\mathbb {P}\left(X_n > x\right) = e^{-\lambda x} and that XnX_n is independent of Sn−1S_{n-1}.

(d)

Argue that XnX_n is independent of X1,X2,…,Xn−1X_1, X_2, \ldots , X_{n-1}.

Problem 2.5
(a)
Problem 2.6

Let h(x)h(x) be a positive function of a real variable that satisfies h(x+t)=h(x)+h(t)h(x+t) = h(x)+h(t) and let h(1)=ch(1) = c.

?
(a)

Show that for integer k>0k > 0, h(k)=kch(k) = kc.

(b)

Show that for integer j>0j > 0, h(1/j)=c/jh(1/j) = c/j.

(c)

Show that for all integer k,jk, j, h(k/j)=ck/jh(k/j) = ck/j.

(d)

The above parts show that h(x)h(x) is linear in positive rational numbers. For very picky mathematicians, this does not guarantee that h(x)h(x) is linear in positive real numbers. Show that if h(x)h(x) is also monotonic in xx, then h(x)h(x) is linear in x>0x > 0.

Problem 2.7

Assume that a counting process {N(t);t>0}\left\{ N(t); t>0\right\} has the independent and stationary increment properties and, for all t>0t > 0, satisfies

P(N~(t,t+δ)=0)=1−λδ+o(δ),P(N~(t,t+δ)=1)=λδ+o(δ),P(N~(t,t+δ)>1)=o(δ). \begin{aligned} \mathbb {P}\left(\widetilde{N}(t, t+\delta ) = 0\right) & = 1-\lambda \delta +o(\delta ), \\ \mathbb {P}\left(\widetilde{N}(t, t+\delta ) = 1\right) & = \lambda \delta +o(\delta ), \\ \mathbb {P}\left(\widetilde{N}(t, t+\delta ) > 1\right) & = o(\delta ). \end{aligned}
?
(a)

Let F0(τ)=P(N(τ)=0)F_0(\tau ) = \mathbb {P}\left(N(\tau ) = 0\right) and show that dF0(τ)/dτ=−λF0(τ)dF_0(\tau )/d\tau = -\lambda F_0(\tau ).

(b)

Show that X1X_1, the time of the first arrival, is exponential with parameter λ\lambda.

(c)

Let Fnc(τ)=P(N~(t,t+τ)=0∣Sn−1=t)F_n^c(\tau ) = \mathbb {P}\left(\widetilde{N}(t, t+\tau ) = 0 \mid S_{n-1} = t\right) and show that dFnc(τ)/dτ=−λFnc(τ)dF_n^c(\tau )/d\tau = -\lambda F_n^c(\tau ).

(d)

Argue that XnX_n is exponential with parameter λ\lambda and independent of earlier arrival times.

Problem 2.8

For a Poisson process, let t>0t > 0 be an arbitrary time and let Z1Z_1 be the duration of the interval from tt until the next arrival after tt. Let ZmZ_m, for each m>1m > 1, be the interarrival time from the epoch of the (m−1)(m-1)th arrival after tt until the mmth arrival.

?
(a)

Given that N(t)=nN(t) = n, explain why Zm=Xm+nZ_m = X_{m+n} for m>1m > 1 and Z1=Xn+1−t+SnZ_1 = X_{n+1}-t+S_n.

(b)

Conditional on N(t)=nN(t) = n and Sn=τS_n = \tau, show that Z1,Z2,…Z_1, Z_2, \ldots are IID.

(c)

Show that Z1,Z2,…Z_1, Z_2, \ldots are IID.

Problem 2.9

Consider a `shrinking Bernoulli' approximation Nδ(mδ)=Y1+⋯+YmN_\delta (m\delta ) = Y_1+\cdots +Y_m to a Poisson process as described in Subsection 2.2.5.

?
(a)

Show that

P(Nδ(mδ)=n)=(mn)(λδ)n(1−λδ)m−n. \mathbb {P}\left(N_\delta (m\delta ) = n\right) = \binom {m}{n} (\lambda \delta )^n (1-\lambda \delta )^{m-n}.
(b)

Let t=mδt = m\delta, and let tt be fixed for the remainder of the exercise. Explain why

lim⁡δ→0P(Nδ(t)=n)=lim⁡m→∞(mn)(λtm)n(1−λtm)m−n, \lim _{\delta \to 0} \mathbb {P}\left(N_\delta (t) = n\right) = \lim _{m \to \infty } \binom {m}{n} \left(\frac{\lambda t}{m}\right)^n \left(1-\frac{\lambda t}{m}\right)^{m-n},

where the limit on the left-hand side is taken over values of δ\delta that divide tt.

(c)

Derive the following two equalities:

lim⁡m→∞(mn)1mn=1n!andlim⁡m→∞(1−λtm)m−n=e−λt. \lim _{m \to \infty } \binom {m}{n} \frac{1}{m^n} = \frac{1}{n!} \qquad \text{and} \qquad \lim _{m \to \infty } \left(1-\frac{\lambda t}{m}\right)^{m-n} = e^{-\lambda t}.
(d)

Conclude from this that for every tt and every nn, lim⁡δ→0P(Nδ(t)=n)=P(N(t)=n)\lim_{\delta \to 0} \mathbb {P}\left(N_\delta (t) = n\right) = \mathbb {P}\left(N(t) = n\right), where {N(t);t>0}\left\{ N(t); t > 0\right\} is a Poisson process of rate λ\lambda.

Problem 2.10

Let {N(t);t>0}\left\{ N(t); t>0\right\} be a Poisson process of rate λ\lambda.

?
(a)

Find the joint probability mass function (PMF) of N(t),N(t+s)N(t), N(t+s) for s>0s > 0.

(b)

Find E[N(t)⋅N(t+s)]\mathbb {E}\left[N(t) \cdot N(t+s)\right] for s>0s > 0.

(c)

Find E[N~(t1,t3)⋅N~(t2,t4)]\mathbb {E}\left[\widetilde{N}(t_1, t_3) \cdot \widetilde{N}(t_2, t_4)\right], where N~(t,τ)\widetilde{N}(t, \tau ) is the number of arrivals in (t,τ](t, \tau ] and t1<t2<t3<t4t_1 < t_2 < t_3 < t_4.

Problem 2.11

An elementary experiment is independently performed NN times, where NN is a Poisson rv of mean λ\lambda. Let {a1,a2,…,aK}\left\{ a_1, a_2, \ldots , a_K\right\} be the set of sample points of the elementary experiment and let pkp_k, 1≤k≤K1 \leq k \leq K, denote the probability of aka_k.

?
(a)

Let NkN_k denote the number of elementary experiments performed for which the output is aka_k. Find the PMF for NkN_k (1≤k≤K1 \leq k \leq K).

(b)

Find the PMF for N1+N2N_1+N_2.

(c)

Find the conditional PMF for N1N_1 given that N=nN = n.

(d)

Find the conditional PMF for N1+N2N_1+N_2 given that N=nN = n.

(e)

Find the conditional PMF for NN given that N1=n1N_1 = n_1.

Remark.
?
(a)

No calculation is necessary.

Problem 2.12

Starting from time 00, northbound buses arrive at 77 Mass. Avenue according to a Poisson process of rate λ\lambda. Customers arrive according to an independent Poisson process of rate μ\mu. When a bus arrives, all waiting customers instantly enter the bus and subsequent customers wait for the next bus.

?
(a)

Find the PMF for the number of customers entering a bus (more specifically, for any given mm, find the PMF for the number of customers entering the mmth bus).

(b)

Find the PMF for the number of customers entering the mmth bus given that the interarrival interval between bus m−1m-1 and bus mm is xx.

(c)

Given that a bus arrives at time 10:30 pm, find the PMF for the number of customers entering the next bus.

(d)

Given that a bus arrives at time 10:30 pm and no bus arrives between 10:30 and 11 pm, find the PMF for the number of customers on the next bus.

(e)

Find the PMF for the number of customers waiting at some given time, say 2:30 pm (assume that the processes started infinitely far in the past).

(f)

Find the PMF for the number of customers getting on the next bus to arrive after 2:30 pm.

(g)

Given that I arrive to wait for a bus at 2:30 pm, find the PMF for the number of customers getting on the next bus.

Remark.
?
(e)

Think of what happens moving backward in time from 2:30 pm.

(f)

This is different from (a); look carefully at part (e).

Problem 2.13
?
(a)

Show that the arrival epochs of a Poisson process satisfy

fS→(n)∣Sn+1(s→(n)∣sn+1)=n!/sn+1n. f_{\overrightarrow {S}^{(n)} \mid S_{n+1}}\left(\overrightarrow {s}^{(n)} \mid s_{n+1}\right) = n!/s_{n+1}^n.
(b)

Contrast this with the result of Theorem 2.5.1.

Remark.
?
(a)

This is easy if you use only the results of Section 2.2.2.

Problem 2.14

Equation (2.42) gives fSi∣N(t)(si∣n)f_{S_i \mid N(t)}(s_i \mid n), which is the density of rv SiS_i conditional on N(t)=nN(t) = n for n≥in \geq i. Multiply this expression by P(N(t)=n)\mathbb {P}\left(N(t) = n\right) and sum over nn to find fSi(si)f_{S_i}(s_i); verify that your answer is indeed the Erlang density.

?
Problem 2.15

Consider generalizing the bulk arrival process in Figure 2.5. Assume that the epochs at which arrivals occur form a Poisson process {N(t);t>0}\left\{ N(t); t>0\right\} of rate λ\lambda. At each arrival epoch SnS_n, the number of arrivals ZnZ_n, satisfies P(Zn=1)=p\mathbb {P}\left(Z_n=1\right) = p, P(Zn=2)=1−p\mathbb {P}\left(Z_n=2\right) = 1-p. The rv's ZnZ_n are IID.

?
(a)

Let {N1(t);t>0}\left\{ N_1(t); t>0\right\} be the counting process of the epochs at which single arrivals occur. Find the PMF of N1(t)N_1(t) as a function of tt. Similarly, let {N2(t);t≥0}\left\{ N_2(t); t \geq 0\right\} be the counting process of the epochs at which double arrivals occur. Find the PMF of N2(t)N_2(t) as a function of tt.

(b)

Let {NB(t);t≥0}\left\{ N_B(t); t \geq 0\right\} be the counting process of the total number of arrivals. Give an expression for the PMF of NB(t)N_B(t) as a function of tt.

Problem 2.16
?
(a)

For a Poisson counting process of rate λ\lambda, find the joint probability density of S1,S2,…,Sn−1S_1, S_2, \ldots , S_{n-1} conditional on Sn=tS_n = t.

(b)

Find P(X1>τ∣Sn=t)\mathbb {P}\left(X_1 > \tau \mid S_n=t\right).

(c)

Find P(Xi>τ∣Sn=t)\mathbb {P}\left(X_i > \tau \mid S_n=t\right) for 1≤i≤n1 \leq i \leq n.

(d)

Find the density fSi∣Sn(si∣t)f_{S_i \mid S_n}(s_i \mid t) for 1≤i≤n−11 \leq i \leq n-1.

(e)

Give an explanation for the striking similarity between the condition N(t)=n−1N(t) = n-1 and the condition Sn=tS_n = t.

Problem 2.17
?
(a)

For a Poisson process of rate λ\lambda, find P(N(t)=n∣S1=τ)\mathbb {P}\left(N(t)=n \mid S_1=\tau \right) for t>τt > \tau and n≥1n \geq 1.

(b)

Using this, find fS1(τ∣N(t)=n)f_{S_1}(\tau \mid N(t)=n).

(c)

Check your answer against (2.41).

Problem 2.18

Consider a counting process in which the rate is a rv Λ\Lambda with probability density fΛ(λ)=αe−αλf_\Lambda (\lambda ) = \alpha e^{-\alpha \lambda } for λ>0\lambda > 0. Conditional on a given sample value λ\lambda for the rate, the counting process is a Poisson process of rate λ\lambda (i.e., nature first chooses a sample value λ\lambda and then generates a sample path of a Poisson process of that rate λ\lambda).

?
(a)

What is P(N(t)=n∣Λ=λ)\mathbb {P}\left(N(t)=n \mid \Lambda =\lambda \right), where N(t)N(t) is the number of arrivals in the interval (0,t](0, t] for some given t>0t > 0?

(b)

Show that P(N(t)=n)\mathbb {P}\left(N(t)=n\right), the unconditional PMF for N(t)N(t), is given by

P(N(t)=n)=αtn(t+α)n+1. \mathbb {P}\left(N(t)=n\right) = \frac{\alpha t^n}{(t+\alpha )^{n+1}}.
(c)

Find fΛ(λ∣N(t)=n)f_\Lambda (\lambda \mid N(t)=n), the density of λ\lambda conditional on N(t)=nN(t)=n.

(d)

Find E[Λ∣N(t)=n]\mathbb {E}\left[\Lambda \mid N(t)=n\right] and interpret your result for very small tt with n=0n = 0 and for very large tt with nn large.

(e)

Find E[Λ∣N(t)=n,S1,S2,…,Sn]\mathbb {E}\left[\Lambda \mid N(t)=n, S_1, S_2, \ldots , S_n\right]. Find E[Λ∣N(t)=n,N(τ)=m]\mathbb {E}\left[\Lambda \mid N(t)=n, N(\tau )=m\right] for some τ<t\tau < t.

Remark.
?
(e)

Consider the distribution of S1,…,SnS_1, \ldots , S_n conditional on N(t)N(t) and Λ\Lambda.

Problem 2.19
?
(a)

Use Equation (2.42) -- fSi∣N(t)(si∣n)=sii−1(i−1)!(t−si)n−i(n−i)!n!tnf_{S_i \mid N(t)}(s_i \mid n) = \frac{s_i^{i-1}}{(i-1)!} \frac{(t-s_i)^{n-i}}{(n-i)!} \frac{n!}{t^n} for n≥in \geq i -- to find E[Si∣N(t)=n]\mathbb {E}\left[S_i \mid N(t)=n\right].

(b)

Find the second moment and the variance of SiS_i conditional on N(t)=nN(t)=n.

(c)

Assume that nn is odd, and consider i=(n+1)/2i = (n+1)/2. What is the relationship between SiS_i, conditional on N(t)=nN(t)=n, and the sample median of nn IID uniform rv's.

(d)

Give a weak law of large numbers (WLLN) for the above median.

Remark.
?
(a)

When you integrate sifSi(si∣N(t)=n)s_i f_{S_i}(s_i \mid N(t)=n), compare this integral with fSi+1(si∣N(t)=n+1)f_{S_{i+1}}(s_i \mid N(t)=n+1) and use the fact that the latter expression is a probability density.

(b)

Extend the previous hint.

Problem 2.20

Suppose cars enter a one-way infinite length, infinite lane highway at a Poisson rate λ\lambda. The iith car to enter chooses a velocity ViV_i and travels at this velocity. Assume that the ViV_i's are independent positive rv's having a common distribution FF. Derive the distribution of the number of cars that are located in an interval (0,a)(0, a) at time tt.

?
Problem 2.21

Consider an M/G/∞\infty queue, i.e., a queue with Poisson arrivals of rate λ\lambda in which each arrival ii, independent of other arrivals, remains in the system for a time XiX_i, where {Xi;i≥1}\left\{ X_i; i \geq 1\right\} is a set of IID rv's with some given CDF F(x)F(x).

You may assume that the number of arrivals in any interval (t,t+ϵ)(t, t+\epsilon ) that are still in the system at some later time τ≥t+ϵ\tau \geq t+\epsilon is statistically independent of the number of arrivals in that same interval (t,t+ϵ)(t, t+\epsilon ) that have departed from the system by time τ\tau.

?
(a)

Let N(τ)N(\tau ) be the number of customers in the system at time τ\tau. Find the mean, m(τ)m(\tau ), of N(τ)N(\tau ) and find P(N(τ)=n)\mathbb {P}\left(N(\tau ) = n\right).

(b)

Let D(τ)D(\tau ) be the number of customers that have departed from the system by time τ\tau. Find the mean, E[D(τ)]\mathbb {E}\left[D(\tau )\right], and find P(D(τ)=d)\mathbb {P}\left(D(\tau ) = d\right).

(c)

Find P(N(τ)=n,D(τ)=d)\mathbb {P}\left(N(\tau ) = n, D(\tau ) = d\right).

(d)

Let A(τ)A(\tau ) be the total number of arrivals up to time τ\tau. Find P(N(τ)=n∣A(τ)=a)\mathbb {P}\left(N(\tau ) = n \mid A(\tau ) = a\right).

(e)

Find P(D(τ+ϵ)−D(τ)=d)\mathbb {P}\left(D(\tau +\epsilon )-D(\tau ) = d\right).

Problem 2.22

The voters in a given town arrive at the place of voting according to a Poisson process of rate λ=100\lambda = 100 voters per hour. The voters independently vote for candidate AA and candidate BB each with probability 1/21/2. Assume that the voting starts at time 00 and continues indefinitely.

?
(a)

Conditional on 1000 voters arriving during the first 10 hours of voting, find the probability that candidate AA receives nn of those votes.

(b)

Again conditional on 1000 voters during the first 10 hours, find the probability that candidate AA receives nn votes in the first 4 hours of voting.

(c)

Let TT be the epoch of the arrival of the first voter voting for candidate AA. Find the density of TT.

(d)

Find the PMF of the number of voters for candidate BB who arrive before the first voter for AA.

(e)

Define the nnth voter as a reversal if the nnth voter votes for a different candidate than the (n−1)(n-1)th. For example, in the sequence of votes AABAABBAABAABB, the third, fourth, and sixth voters are reversals; the third and sixth are AA to BB reversals and the fourth is a BB to AA reversal. Let N(t)N(t) be the number of reversals up to time tt (tt in hours). Is {N(t);t>0}\left\{ N(t); t > 0\right\} a Poisson process? Explain.

(f)

Find the expected time (in hours) between reversals.

(g)

Find the probability density of the time between reversals.

(h)

Find the density of the time from one AA to BB reversal to the next AA to BB reversal.

Problem 2.23

Let {N1(t);t>0}\left\{ N_1(t); t > 0\right\} be a Poisson counting process of rate λ\lambda. Assume that the arrivals from this process are switched on and off by arrivals from a second independent Poisson process {N2(t);t>0}\left\{ N_2(t); t > 0\right\} of rate γ\gamma.

 {#fig-1 width="85%"}

Let {NA(t);t≥0}\left\{ N_A(t); t \geq 0\right\} be the switched process, i.e., NA(t)N_A(t) includes the arrivals from {N1(t);t>0}\left\{ N_1(t); t > 0\right\} during periods when N2(t)N_2(t) is even and excludes the arrivals from {N1(t);t>0}\left\{ N_1(t); t > 0\right\} while N2(t)N_2(t) is odd.

?
(a)

Find the PMF for the number of arrivals of the first process, {N1(t);t>0}\left\{ N_1(t); t > 0\right\}, during the nnth period when the switch is on.

(b)

Given that the first arrival for the second process occurs at epoch τ\tau, find the conditional PMF for the number of arrivals of the first process up to τ\tau.

(c)

Given that the number of arrivals of the first process, up to the first arrival for the second process, is nn, find the density for the epoch of the first arrival from the second process.

(d)

Find the density of the interarrival time for {NA(t);t≥0}\left\{ N_A(t); t \geq 0\right\}.

Remark.
?
(d)

This part is quite messy and is done most easily via Laplace transforms.

Problem 2.24

Let us model the chess tournament between Fisher and Spassky as a stochastic process. Let XiX_i, for i≥1i \geq 1, be the duration of the iith game and assume that {Xi;i≥1}\left\{ X_i; i \geq 1\right\} is a set of IID exponentially distributed rv's each with density fX(x)=λexp(−λx)f_X(x) = \lambda \mathrm{exp}\left(-\lambda x\right). Suppose that each game (independently of all other games, and independently of the length of the games) is won by Fisher with probability pp, by Spassky with probability qq, and is a draw with probability 1−p−q1-p-q. The first player to win nn games is defined to be the winner, but we consider the match up to the point of winning as being embedded in an unending sequence of games.

?
(a)

Find the distribution of time, from the beginning of the match, until the completion of the first game that is won (i.e., that is not a draw). Characterize the process of the number {N(t);t>0}\left\{ N(t); t > 0\right\} of games won up to and including time tt. Characterize the process of the number {NF(t);t≥0}\left\{ N_F(t); t \geq 0\right\} of games won by Fisher and the number {NS(t);t≥0}\left\{ N_S(t); t \geq 0\right\} won by Spassky.

(b)

For the remainder of the problem, assume that the probability of a draw is zero, i.e., that p+q=1p+q = 1. How many of the first 2n−12n-1 games must be won by Fisher in order to win the match?

(c)

What is the probability that Fisher wins the match? Your answer should not involve any integrals.

(d)

Let TT be the epoch at which the match is completed (i.e., either Fisher or Spassky wins). Find the CDF of TT.

(e)

Find the probability that Fisher wins and that TT lies in the interval (t,t+δ)(t, t+\delta ) for arbitrarily small δ\delta.

Remark.
?
(c)

Consider the unending sequence of games and use (b).

Problem 2.25
?
(a)

For 1≤i<n1 \leq i < n, find the conditional density of Si+1S_{i+1}, conditional on N(t)=nN(t) = n and Si=siS_i = s_i.

(b)

Use (a) to find the joint density of S1,…,SnS_1, \ldots , S_n conditional on N(t)=nN(t) = n. Verify that your answer agrees with (2.38).

Problem 2.26

A two-dimensional Poisson process is a process of randomly occurring special points in the plane such that (i) for any region of area AA the number of special points in that region has a Poisson distribution with mean λA\lambda A, and (ii) the number of special points in non-overlapping regions is independent. For such a process consider an arbitrary location in the plane and let XX denote its distance from its nearest special point (where distance is measured in the usual Euclidean manner). Show that:

?
(a)

P(X>t)=exp(−λπt2)\mathbb {P}\left(X > t\right) = \mathrm{exp}\left(-\lambda \pi t^2\right);

(b)

E[X]=1/(2λ)\mathbb {E}\left[X\right] = 1/(2\sqrt{\lambda }).

Problem 2.27

This problem is intended to show that one can analyze the long term behavior of queueing problems by using just notions of means and variances, but that such analysis is awkward, justifying understanding the strong law of large numbers (SLLN). Consider an M/G/1 queue. The arrival process is Poisson with λ=1\lambda = 1. The expected service time, E[Y]\mathbb {E}\left[Y\right], is 1/21/2 and the variance of the service time is 1.

?
(a)

Consider SnS_n, the time of the nnth arrival, for n=1012n = 10^{12}. With high probability, SnS_n will lie within three standard deviations of its mean. Find and compare this mean and the 3σ3\sigma range.

(b)

Let VnV_n be the total amount of time during which the server is busy with these nn arrivals (i.e., the sum of 101210^{12} service times). Find the mean and 3σ3\sigma range of VnV_n.

(c)

Find the mean and 3σ3\sigma range of InI_n, the total amount of time the server is idle up until SnS_n (take InI_n as Sn−VnS_n-V_n, thus ignoring any service time after SnS_n).

(d)

An idle period starts when the server completes a service and there are no waiting arrivals; it ends on the next arrival. Find the mean and variance of an idle period. Are successive idle periods IID?

(e)

Combine (c) and (d) to estimate the total number of idle periods up to time SnS_n. Use this to estimate the total number of busy periods.

(f)

Combine (e) and (b) to estimate the expected length of a busy period.

Problem 2.28

The purpose of this problem is to illustrate that for an arrival process with independent but not identically distributed interarrival intervals, X1,X2,…X_1, X_2, \ldots, the number of arrivals N(t)N(t) in the interval (0,t](0, t] can be a defective rv. In other words, the `counting process' is not necessarily a stochastic process according to our definitions. This also suggests that it might be necessary to prove that the counting rv's for a renewal process are actually rv's.

?
(a)

Let the CDF of the iith interarrival interval for an arrival process be FXi(xi)=1−exp(−α−ixi)F_{X_i}(x_i) = 1-\mathrm{exp}\left(-\alpha^{-i} x_i\right) for some fixed α∈(0,1)\alpha \in (0,1). Let Sn=X1+⋯+XnS_n = X_1+\cdots +X_n and show that

E[Sn]=α(1−αn)1−α. \mathbb {E}\left[S_n\right] = \frac{\alpha (1-\alpha ^n)}{1-\alpha }.
(b)

Sketch a `reasonable' sample path for N(t)N(t).

(c)

Use the Markov inequality on P(Sn≥t)\mathbb {P}\left(S_n \geq t\right) to find an upper bound on P(N(t)<n)\mathbb {P}\left(N(t) < n\right) that is smaller than 1 for all nn and for large enough tt. Use this to show that N(t)N(t) is defective for large enough tt.

(d)

(For those looking for a challenge) Show that N(t)N(t) is defective for all t>0t > 0.

Remark.
?
(d)

Use the Markov inequality to find an upper bound on P(Sm−Sn≤t/2)\mathbb {P}\left(S_m-S_n \leq t/2\right) for all m>nm > n for any fixed nn. Show that, for any t>0t > 0, this is bounded below 1 for large enough nn. Then show that SnS_n has a density that is positive for all t>0t > 0.